Question code 65 56 70 Stockholms universitet .A zero-coupon bond with maturity in two years and face value 1000 SEK has a price equal to 990 SEK. What is its yield-to-maturity? A. 1.01% B. 2.02% -: r v C. 0.50% D. 3.00% E. It cannot be computed FV (I+YTM)M
Answer #1

P = FV / ( 1 + YTM)n

P / FV = 1 / ( 1 + YTM)n

990 / 1000 = 1 / ( 1 + YTM)2

0.99 = 1 / ( 1 + YTM)2

( 1 + YTM)2 = 1 / 0.99

( 1 + YTM)2 = 1.010101

1 + YTM = 1.0050

YTM = 0.0050 or 0.50%